what is the lenght of the hypotenus of the right triangle.show the work on a graph.a is 4, b-5 what is c
Answer by nyc_function(2741)

a^2 + b^2 = c^2
4^2 + 5^2 = c^2
16 + 25 = c^2
41 = c^2
sqrt{41} = sqrt{c^2}
sqrt{41} = c
Done!